40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 29 19:38:09 2015, gautam, Update, General, FOL fiber box revamp IMG_0017.JPGFOL_schematic.pdfIMG_0018.JPGIMG_0016.JPG
    Reply  Wed Sep 30 13:07:13 2015, gautam, Update, General, FOL fiber box revamp FOL_schematic.pdf
       Reply  Tue Oct 6 16:56:40 2015, gautam, Update, General, FOL fiber box revamp 20151006_Xbeat_psd.pdf20151006_Xbeat_coherence.pdf
Message ID: 11670     Entry time: Tue Oct 6 16:56:40 2015     In reply to: 11652
Author: gautam 
Type: Update 
Category: General 
Subject: FOL fiber box revamp 

[gautam, ericq]

We had a look at the IR beat (PSL+Xarm) today using the new FOL fiber box, and compared it to the green beat signal for the same combination. We first switched out the green Y beat input into the RF amplifiers on the PSL table with the PSL+Xarm IR beat input (so in all the plots, the BEATY channels really correspond to the IR beat for PSL+X). The IR and green beat notes were found without much difficulty, and we compared the beat signal PSDs for the green and IR signals (see Attachment #1 - arms were locked to green and the X slow control was turned on). The pink trace (labeled REF1) corresponds to the green beat signal, and was in good agreement with an earlier reference trace Eric had saved for the same signal. The teal trace (labeled REF0) corresponds to the the IR beat signal monitored simultaneously. 

We then went back to the PSL table to check the amplitude of the signal from the broadband fiber PDs using the Agilent network analyzer. An initial measurement yielded a beat note (@~50MHz) at ~-22dbm (17mV rms). We figured that by bypassing the 90-10 splitter in this path, we could get a stronger signal. But after switching out the fiber connections we found that the signal amplitude had fallen to ~-27dbm (10mV rms). As per my earlier measurements here, we expect ~600uW of light on the PD, and a quick calculation suggested the signal should be more like 60mV, so we used the fiber power meter to check the power levels after each of the couplers again. We then found that the fiber connector on the front panel of the box for the PSL input wasnt ideal (the laser power after the first 50-50 coupler was only ~250 uW, though the input was ~1.2  mW). The power after the first coupler also fluctuated unpredictably (<100 uW to 350 uW) in response to slightly tightening/loosening the fiber connections on the front panel. I then switched the PSL input to one of the two unused fiber connectors on the front panel (meant for the 10% of the beat signal for the DC readout), and found that this input behaved much better, with ~450 uW of power available after the first 50-50 coupler. The power going into the beat PD was also measured to be ~550uW, closer to what was expected. The beat signal peak now was ~-14dbm (~30mV rms).

We then once again repeated the comparison between green and IR beat signals - but while in the control room, I noticed that the beat signal amplitude on the network analyzer in the control room was fluctuating by nearly 1.5 divisions on the vertical scale - not sure what the reason for this is. A look at the PSD of the IR beat with higher power incident on the PD was also not encouraging (see blue trace in Attachment #1), it seems to have gotten worse in the 10-30 Hz range. We also looked at the coherence between the beat spectrum and the beat note amplitude in order to look for any linear coupling between the two, but from Attachment #2, we cannot explain the disparity between the green and IR beat spectra. This warrants further investigation.

Everything on the PSL table has now been restored to the configurations before these investigations (i.e. the Y+PSL green beat cable has been reconnected to the RF amplifier, and both green beat PDs have been powered back ON. The fiber PDs are powered OFF) 

Attachment 1: 20151006_Xbeat_psd.pdf  50 kB  | Hide | Hide all
20151006_Xbeat_psd.pdf
Attachment 2: 20151006_Xbeat_coherence.pdf  44 kB  | Hide | Hide all
20151006_Xbeat_coherence.pdf
ELOG V3.1.3-